OpenStudy (anonymous):

As VIN is increased, the output voltage VOUT decreases and the MOSFET goes out of saturation. For the value or RL found above, what is the maximum input voltage VIN in volts such that the MOSFET will remain in the saturation region?

OpenStudy (anonymous):

8k and -0.5

OpenStudy (anonymous):

What is the maximum input voltage in Vin?

OpenStudy (anonymous):

for these values .. VS+=1.0V, VS−=−1.0V, and the MOSFET parameters are K=1 mAV2 and VT=0.5V. the answer is 0.093

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

I got it!!! Actually in finding valid Operating ranges \[V0= [ (-1+\sqrt{1+2*K*RL*Vs}) ]\div KRL\] In the above formula Vs= applied voltage across RL and Drain to Source( i.e Vds).. \[Vs= ids*RL + Vds (or) Vs=ids*RL + V0\] But in given problem , by applying KVL to output loop we get \[Vds+idsRL = Vs+ - Vs-\] \[Vds+ids*RL = 1 - (-1) =2 \] So in finding V0 substitute Vs=2 (NOT 1)... And we know Valid maximum input range is Vgs = V0 + Vthershold Applying KVL to input loop \[Vin = Vgs + Vs-\] Substitue the value ...You will get it !!!!!! Cheers!!!!

OpenStudy (anonymous):

how will u find RL?

OpenStudy (anonymous):

\[Vin=Vgs+(Vs-)\]\[Vout=(Vs+) -ids*RL\]\[ids=(Vgs-Vt)^2 *K/2\]Now you have all the required Vin and Vout , So go and subtitute to find RL... Cheers!!